xy (z2 + 1) + z (x2 + y2) can be factorized as:


(A) (xy + z) (yz + x) (B ) (zx + y) (xy + z)


(C) (yz + x ) (zx + y) (D) (x2 + y2) (z + x)

Grade:Upto college level

1 Answers

Pratik Tibrewal
10 years ago
it is (C)
xyz^2 + xy + zx^2 + zy^2
= xyz^2 + zx^2 + xy + zy^2
= xz(yz + x) +y(zy + x)
= (yz + x)(xz + y)


Thanks and Regards
Pratik Tibrewal,
askiitians faculty,
BTech